A stone is thrown horizontal with an initial velocity of 20m/s from a top of clift with a height of 60m. Determine the range


1
Expert's answer
2021-04-19T07:24:30-0400

Answer

Range is given by

"R=u\sqrt{\frac{2H}{g}}\\=20\times\sqrt{\frac{2\times60}{9.8}}"

=71.1m